40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 9 15:31:45 2017, gautam, Update, IMC, IMC wonkiness IMG_7409.JPG
    Reply  Fri Jun 9 17:45:21 2017, Gautam, Kaustubh, Update, IMC, IMC wonkiness 
       Reply  Fri Jun 9 19:18:10 2017, gautam, Update, IMC, IMC wonkiness MC2_UL_glitchy.pngMC2_glitch_fast.png
          Reply  Mon Jun 12 22:23:20 2017, rana, Update, IMC, IMC wonkiness 
          Reply  Tue Jun 13 08:40:32 2017, Steve, Update, IMC, IMC wonkiness happy_MC.pnglast_glitch.png
Message ID: 13062     Entry time: Tue Jun 13 08:40:32 2017     In reply to: 13058
Author: Steve 
Type: Update 
Category: IMC 
Subject: IMC wonkiness 

Happy MC after last glitch at 10:28 so the credit goes to Rana

GV edit 11:30am: I think the stuff at 10:28 is not a glitch but just the WFS servos coming on - the IMC was only hand aligned before this.

Quote:

It happened again. MC2 UL seems to have gotten the biggest glitch. It's a rather small jump in the signal level compared to what I have seen in the recent past in connection with suspect Satellite boxes, and LL and UR sensors barely see it.

I will squish Sat box cables and check the cabling at the coil driver board end as well, given that these are two areas where there has been some work recently. WFS loops will remain off till I figure this out. At least the (newly centered) DC spot positions on the WFS and MC2 TRANS QPD should serve as some kind of reference for good MC alignment.

GV edit 9pm: I tightened up all the cables, but doesn't seem to have helped. There was another, larger glitch just now. UR and LL basically don't see it at all (see Attachment #2). It also seems to be a much slower process than the glitches seen on MC1, with the misalignment happening over a few seconds (it is also a lot slower). I have to see if this is consistent with a glitch in the bias voltage to one of the coils which gets low passed by a 4xpole@1Hz filter.

Quote:

Once we had the beam approximately centered for all of the above 3 PDs, we turned on the locking for IMC, and it seems to work just fine. We are waiting for another hour for switching on the angular allignment for the mirrors to make sure the alignment holds with WFS turned off.

 

 

Attachment 1: happy_MC.png  1.051 MB  | Hide | Hide all
happy_MC.png
Attachment 2: last_glitch.png  44 kB  Uploaded Tue Jun 13 10:09:48 2017  | Hide | Hide all
last_glitch.png
ELOG V3.1.3-